### v1

# SAVeD: Learning to Denoise Low-SNR Video for Improved Downstream Performance

Abstract
Foundation models excel at vision tasks in natural images but fail in low signal-to-noise ratio (SNR) videos, such as underwater sonar, ultrasound, and microscopy. We introduce Spatiotemporal Augmentations and denoising in Video for Downstream Tasks (SAVeD), a self-supervised method that denoises low-SNR sensor videos and is trained using only the raw noisy data. By leveraging differences in foreground and background motion, SAVeD enhances object visibility using an encoder-decoder with a temporal bottleneck. Our approach improves classification, detection, tracking, and counting, outperforming state-of-the-art video denoising methods with lower resource requirements. Project page: https://suzanne-stathatos.github.io/SAVeD/. Code page: https://github.com/suzanne-stathatos/SAVeD.
1 Introduction
Motion may be the only way to identify objects in video with low signal-to-noise-ratio (SNR), camouflage, or complex textures that may hinder frame-by-frame object detection. The human visual system is excellent at capturing observable motion [26], and this capability has not yet been reproduced by modern generative models. Learning to exploit motion cues will improve models’ ability to detect and track objects of interest in noisy video.
Obtaining sufficient annotations to train supervised models in video can be prohibitively expensive, especially for scientific [30] or medical [64, 15] applications, and delays the deployment of models when tackling novel signal statistics. Once trained, supervised detectors may not generalize well and may need additional annotations to be adapted to other locations with different background and foreground distributions[49], which may require additional supervision [14]. On the other hand, models trained using self-supervision can be more robust [60, 47]. This work aims to enhance motion signals in unlabeled low-SNR data, such as ultrasound and sonar videos, to improve downstream supervised classification, detection, and tracking.
Figure 1: SpatioTemporal Denoising improves classification, detection, tracking, and counting in video. We denoise sonar and ultrasound videos of fish in a river, lung scans, breast lesion scans, and cell microscopy to improve downstream classification, detection, tracking, and counting tasks. We propose a self-supervised method to enhance the foreground signal of video frames without manual annotations. Our method works on videos with: non-stationary backgrounds, low signal-to-noise-ratios, and a variable number of objects in a video.
Unsupervised and self-supervised methods are increasingly used for object localization and action recognition [38, 62, 16, 68, 58, 63]. However, existing methods do not address low-SNR videos. Furthermore, while some methods [21, 24] handle changes in camera view-angle, they do not handle cases where the camera is stationary and the background is not.
We address these challenges with SAVeD, a self-supervised learning method to denoise video. Additionally, we exploit object motion to boost the SNR across frames. Inspired by work on self-supervised reconstruction [38, 62], broad vs. narrow self-supervised video understanding [53], and anomaly detection [77, 45], we use an encoder to encode appearance frames, an hourglass network to combine temporal features, and a decoder network to reconstruct the denoised frame.
Our main contributions are:
  • We propose SAVeD, a novel denoising approach to clarify low-SNR video with variable numbers of agents; details are in Sec. 3.
  • We propose a rich benchmark for low-SNR video denoising consisting of a diverse collection of low-SNR video domains (sonar video of fish, ultrasound video of lungs and breast, microscopy video of tissue) and a diverse collection of downstream visual tasks (classification, detection, tracking, and counting), in Sec. 4.
  • We explore the value of different algorithmic choices in low-SNR video processing and test SAVeD, together with a number of variants (in Sec. 5 & A, and Supplementary materials), on our benchmark.
Figure 2: SAVeD, our approach for self-supervised denoising using spatiotemporal difference and identity reconstruction. $I_{t}$, $I_{t-T}$, and $I_{t-2T}$ are video frames at times t (current frame), t-T, and t-2T. These frames are input to an appearance encoder $\Phi$. The resulting feature representations are input to a spatiotemporal bottleneck $\Theta$ that compresses the 3 appearance features into a single spatiotemporal feature representation. Our model then predicts the reconstruction target, defined in Eq. 2 in Sec. 3.2, using the reconstruction decoder $\Psi$. The architecture is discussed in more detail in Sec. 3.3.
2 Related Work
Sonar and Ultrasound Sonar and ultrasound pose a distinct and interesting set of challenges to the computer vision community. Particularly noteworthy elements include: all sonar has pink noise in it, all non-cavity objects of interest have higher pixel intensity values compared to the background, and objects of interest may not be uniquely identifiable from their appearance features. Weld et al. [72] attempt to standardize ultrasound data through geometric analysis and augmentation, given ultrasound-data’s sensor variability. Unlike most computer vision datasets, which focus on light-based data, ultrasound, sonar, lidar, and radar do not rely on capturing light intensity. Instead, they rely on the principle of emitting waves, which bounce off objects, and return to the sensor as echoes. The “camera”, then, measures the distance to those objects by calculating the time it took for the echo to return. Sonar and ultrasound use sound waves, while lidar uses laser light pulses; sonar and ultrasound are primarily used in liquids while lidar is more often used for land and air-based applications [25, 9, 10]. We focus on one sonar dataset and two ultrasound datasets described in more detail in Sec. 4.1.
Classical image denoising. A large number of spatial filters from image processing techniques have been applied to image denoising [73, 67, 51, 4, 75, 65, 23] – they can be broken down to two types: linear and non-linear filters. Linear spatial filters. Mean and Gaussian filtering [27] reduces Gaussian noise, however, they can over-smooth noisy images [1]. Weiner filtering [4] aims to overcome this drawback, but it can overly blur sharp edges. Non-linear filters. With non-linear filters, noise can be reduced without first identifying the noisy pixels. Median-filtering [27] replaces each pixel with the median value of its neighboring pixels. Median and bilateral filtering [67] preserves edges while smoothing images to reduce the noise, though bilateral filtering is inefficient [23].
Self-supervised and unsupervised image denoising. Several approaches use variants of blind-spot networks or pixel-wise masking to denoise imagery. Noise2Self [3] and Noise2Void (N2V) [40] train on noisy images without requiring clean targets or paired noisy data. N2V trains a blind-spot network to predict masked pixels’ intensity values based on neighboring pixels. Others [42, 3, 33] refrain from masking the pixels via a structural blind-spot network composed of half-plane receptive-field U-Nets [55]. Jang et al. [33] use a conditional blind-spot network and a loss that regularizes the denoised images without masking input pixels to train their network. Neighbor2Neighbor [31] proposes a self-supervised loss between two sub-sampled images. In general, noise in real-world imagery, including acoustic imagery, has unknown or non-stationary statistics that are spatially correlated, violating assumptions of pixel-wise independence.
CNN-based video denoising. Some video denoising methods leverage videos’ spatio-temporal structure by using optical flow for motion compensation [74, 66]. DVDnet [66] uses calculated flow-estimates to manually warp frames, align their contents, and process them collectively with a CNN. UDVD [59] uses a patch-wise noise-to-noise training strategy to predict clean frames by estimating masked pixels from adjacent neighborhoods of noisy frames.
Denoising autoencoders (DAEs) were originally introduced to learn more robust representations. During training, DAEs intentionally add noise to their input data and learn to reconstruct the original uncorrupted signal. mDAE [20], a method for missing data imputation (replacing missing or unavailable data), improves performance on a handful of datasets. Zaki et al. [78] leverages a DAE framework as a preprocesing step to improve the quality of SERS spectra for biomarker quantification and discovery. They generated noisy data by duplicating background measurements at random locations. DAEs have also increasingly been applied to video tasks. CompDAE [50] explicitly models noise from snapshot compressive imaging measurements in low-light conditions to improve edge detection and depth estimation. TADA [12] uses an adversarial denoising autoencoder to remove EMG noise from EEG time series data. Our work similarly extends the application of DAEs to spatio-temporal sonar and ultrasound video denoising; we uniquely combine temporal frames to enhance signal quality while simultaneously addressing the increased noise introduced by this process.
Detection by tracking (DbT) methods locate objects in the first frame of a video and then track them to predict future locations [52, 2, 76, 48]. Point-tracking is similar, in that pixels are first initialized and then tracking models are trained to follow the pixels in a video [79, 17, 18, 70, 19, 39, 36, 35]. If the tracking is robust, detection is solved, even in frames where a detector alone fails. Tracking, therefore, helps fill the gap where detectors struggle. DbT is used in video object segmentation [52] to propagate masks over time, in long-term object tracking [34] which re-detects objects when needed, and in Visual Simultaneous Localization and Mapping, which relies on tracking features across frames to infer their presence or position of objects or the camera [11]. In an underwater setting, point-DbT for object discovery or detection may struggle due to the turbidity of water – discovering erraneous points since the background has motion. Inspired by DbT approaches, our method adds temporal information directly to the spatial dimension to boost the foreground signal.
Tracking by detection (TbD) [28, 13] approaches divide the tracking problem into two steps: first, an object detector predicts objects and their locations in every frame; then, a tracker associates detections from one frame to another, creating trajectories and uniquely identifying objects. While there has been notable progress in TbD recently [13], there are also particular drawbacks. TbD relies heavily on detection performance – false alarms, detection gaps, and missed object groupings may lead to incomplete tracks. In addition, some recent progress in TbD can be attributed to matching targets with spatially re-identifiable features. However, this is unrealiable in sonar. Kay et al. [37]’s downstream tracking is a tracking-by-detection approach; in order to improve it, then, we design our method to maximize downstream detection performance by lowering false positive and false negative rates concurrently.
3 Method
The goal of SAVeD (Fig. 2) is to remove noise while capturing and focusing on motion of objects of interest from video with a non-stationary, fluid background. Inspired by previous methods [62, 32, 57], we use an encoder-decoder setup. We propose a novel reconstruction target based on spatiotemporal differences in a neighborhood of frames. We rely on a background with a spatiotemporal distribution that is distinct from the foreground objects.
3.1 Self-supervised denoising
In low-SNR videos, signals are often distributed across multiple frames; as such, we want to condense information from multiple times into a single frame to exaggerate the signal. We do this through the reconstruction target. For simplicity, we choose to reconstruct the spatiotemporal combination of 3 frames, the current input frame $I_{t}$, the future frame $I_{t+T}$ and the previous frame $I_{t-T}$, from three input frames, $I_{t}$, $I_{t-T}$, and $I_{t-2T}$. We explore a vanilla autoencoder, UNets, and 3D convolutions (in Tab. 2), but ultimately find an encoder, bottleneck, and decoder framework works optimally.
We use an encoder-decoder architecture, seen in Fig. 2, with a spatial encoder, $\Phi$, a temporal hourglass network, $\Theta$, and a reconstruction decoder, $\Psi$. During training, spatial encoders, $\Phi$, takes $I_{t}$, $I_{t-T}$, and $I_{t-2T}$ as input to generate spatial feature embeddings, which are then used by the hourglass network, $\Theta$, to generate a spatiotemporal feature embedding; this embedding passes through $\Psi$ to reconstruct the learning objective $\hat{S}_{t,T}$.
$\hat{S}_{t,T}=\Psi(\Theta(\text{concat}(\Phi(I_{t}),\Phi(I_{t-T}),\Phi(I_{t-2T}))))$ (1)
3.2 Reconstruction target and loss
Target. We use the positive frame difference with the current frame (PFDwTN), which incorporates spatiotemporal information as our main reconstruction target.
This combines the current frame with the positive motion from the previous and next frames. Positive motion of the next frame is defined as $\max(0,I_{t}-I_{t+T})$, while positive motion from the previous frame is defined as $\max(0,I_{t}-I_{t-T})$. Note that the previous frame $I_{t-T}$ goes into the network, whereas the future frame $I_{t+T}$ does not. It is seen only when calculating the ground-truth target.
To handle frames where the background movement does not differ significantly from the foreground objects’ motion (i.e., stationary objects), we include the original frame, $I_{t}$, in the reconstruction target. These signals boost the spatial signal by exploiting the motion signature. The overall target is:
$S_{t,T}=\max(0,I_{t}-I_{t-T})+I_{t}+\max(0,I_{t}-I_{t+T})$ (2)
Other motion-augmenting targets that we tested are defined and visualized in Sec. A.1 and Fig. 9 in the Supplemental materials.
Loss. We apply mean-squared-error loss for reconstructing the current frame with augmented motion signatures:
$\mathcal{L}_{recon}=\|\hat{S}_{t,T}-S_{t,T}\|_{2}$ (3)
3.3 Noise Removal Network
Appearance Encoder $\Phi$. We implement a 6-layer CNN. Each layer consists of a convolutional block (Conv2D + ReLU) followed by max pooling, progressively increasing the number of feature channels while reducing the spatial dimension, $\mathbb{R}^{(H,W,1)}\rightarrow\mathbb{R}^{(\frac{H}{32},\frac{W}{32},512)}$. We also save skip connections, which are sequential max pools followed by 1x1 convolutions, to be used by the hourglass network and decoder. This design mimics the encoding portion of a UNet, with a fraction of the parameters and FLOPS, to let the network capture multi-scale features efficiently.
Temporal Hourglass Network $\Theta$ is an hourglass network with a bottleneck consisting of two 3x3 convolutional layers with 512 channels, each followed by ReLU activation. We also have skip connections as feature combiners at each level of the network, designed to merge information from the provided appearance features’ skip connections.
Reconstruction Decoder $\psi$ has 6 upsampling stages, each consisting of a ConvTranspose followed by convolutions and ReLU activations. At each layer, the skip connections from the corresponding encoder level are concatenated with the upsampled features. The decoder reduces the number of channels while increasing the spatial dimension ending with a single-channel output, $\mathbb{R}^{(\frac{H}{32},\frac{W}{32},512)}\rightarrow\mathbb{R}^{(H,W,1)}$.
More details can be seen on each of these in the Supplemental materials Tab. 6.
We recognize that combining noisy frames adds to the noise of the overall signal rather than removing it. Work [5] has shown that denoising methods capture clean data’s underlying structure. Denoising autoencoders purposefully corrupt input data by adding noise or masking some of the input values [68, 29]. We rely on the autoencoder to remove noise implicitly by focusing on the largest reconstruction areas to minimize loss. This assumes that the objects of interest are larger than the noise signature.
3.4 Denoising Metric
Typically [40, 42, 33, 31, 59], denoising networks use Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) [71] as evaluation metrics. PSNR is the ratio of maximum signal power to noise power, and SSIM measures perceived image quality. Both metrics rely on having clean imagery to compare with. Our denoising approach is unsupervised and we do not have clean imagery. As a result, we design a pseudo-PSNR metric to indicate denoised performance for object detection, and we rely on downstream performance for other tasks.
3.4.1 Pseudo-PSNR for Downstream Detection
Recall that we assume that our downstream models are supervised. Therefore, we can assume we have bounding boxes or segmentation masks for detection tasks. For simplicity, we call detection annotations “boxes”, though the same approach works for segmentation masks. For each object’s box $b$ in each frame $I$ of each video, we take the density of pixel intensity values: $d_{b}=I[b]$, where $b$ are the indices associated with the box. Then, we take the density of pixel intensity values from a different frame $\tilde{I}$ of the same video at the same box location $b$ where we know there is no object, $\tilde{d}_{b}$. These densities give us pseudo-distributions between objects (signals) and background (noise). We postulate that if the distributions are separable i.e., the distribution of object pixels is distinct from the distribution of background pixels, then the denoising method works as intended. Therefore, we calculate the distance between object and non-object via the Kullback-Leibler (KL) Divergence [41]. KL divergence measures the distance between two probability distributions P and Q as follows:
[…]
To generate a metric for a data split, we average the $D_{KL}$ over all N bounding boxes to get
$PSNR_{D_{KL}}=\frac{1}{N}\sum_{b\in{N}}D_{KL}(d_{b}\|\tilde{d}_{b})$ (5)
A visualization of this metric can be seen in Supplementary materials Fig. 11.
4 Experiments
We demonstrate that SAVeD can improve performance in low-SNR videos across medical and ecological applications (Sec. 5). We evaluate our denoised images on downstream tasks for detection, tracking, counting, and classification.
4.1 Datasets
Caltech Fish Counting 2022 (CFC22) [37] is designed for detection, tracking, and counting fish in low-signal-to-noise sonar video. This dataset contains 1,567 sonar videos from seven different cameras on three rivers in Alaska and Washington. The videos are grayscale, their resolutions range from 288x624 to 1,086x2,125, their frame rates range from 6.7 to 13.3 fps, and each video is on average 336 frames (38s) in duration [37]. In total, there are 527,215 frames with 8,254 unique fish, totaling 516k bounding boxes and 16.7 hours of video [37]. The dataset includes significant domain shifts (e.g., background topology, occlusion, fish densities, fish sizes, camera noise), requiring models to generalize effectively across varying conditions.
Breast Lesion Ultrasound Video Dataset [44] (BUV) is designed for detection and classification (benign or malignant) of breast lesions. The dataset contains 188 videos, of which 113 are malignant and 75 are benign. These videos collectively have 25,272 images, each with 1 detection; the number of ultrasound images in each video range from 28 to 413. Each video has a complete scan of the abnormal tissue. The dataset has a random train–test split of 150–38 videos respectively[44].
The Point-of-care Ultrasound dataset (POCUS) [8, 7] is a collection of convex and linear probe lung ultrasound images and videos to classify/diagnose COVID-19 and pneumonia. It contains 247 videos and 59 images from both convex and linear probes. We exclusively use the video portion of this set. There are 70 ultrasound videos showing COVID cases, 45 showing possible COVID, 51 videos of bacterial pneumonia, 6 videos of viral pneumonia, and 75 of healthy lungs. Videos are sampled at 10Hz (10 frames per second). We group frames by video as in Born et al. [8, 7]. In total, we extract 9,184 frames. The average width x height of the frames is 499 x 463 pixels.
Fluorescence microscopy dataset [69] (Fluo) is a dataset of fluorescence-microscopy recordings of live cells in [69]. We use the same videos as UDVD [59]: Fluo-32DL-MSC (CTC-MSC), of mesenchymal stem cells, and Fluo-N2DH-GOWT1 (CTC-N2DH), of GOWT1 cells. This dataset also contains no ground-truth clean data. There are a total of 560 frames and four videos.
4.2 Training Procedure
We train SAVeD using the reconstruction objective in Section 3. During training, we rescale CFC22 and POCUS images to 1024x512 and BUV and Fluo images to 1024x1024. For POCUS, we use T of 0.1 seconds (10Hz), as that is what the downstream process uses. For all other datasets, we use all frames. For each dataset, we train over all splits. We train for 20 epochs for CFC22, 120 epochs for POCUS, 40 epochs for BUV, and 1000 epochs for Fluo; we found these numbers of epochs sufficient for training to converge. These took 20 hours, 0.5 hours, 2 hours, and 2 hours, respectively, on 2 RTX 4090 GPUs; this is less time than other network-based denoising methods as seen in Tab. 4 in the Supp Mat. Additional details, including hyperparameter configurations, are in the Supp Mat Sec. B.2.
After training SAVeD, we generate denoised frames for all splits. In the case of CFC22, we combine the denoised image as two channels and the background-subtracted frame, $(I_{v})_{t}-\bar{I_{v}}$, as the last channel. For POCUS, BUV, and Fluo, we combine the the denoised image as two channels and the median-filtered image as the last channel.
Figure 3: Qualitative raw-denoised pairs of SAVeD. Qualitative results for SAVeD trained on POCUS (lung health categorization), BUV (breast lesion detection), CFC22 (fish detection, tracking, and counting), and Fluo (cell denoising).
4.3 Evaluation procedure
Given that none of our videos have clean (noise-free) versions, we use the downstream performance tasks’ metrics as proxies for our denoised performance. We also use our metric from Sec. 3.4.
Denoising for Detection, Tracking, and Counting. For CFC22, which has detection, tracking, and counting as downstream tasks, we follow a simplified version of the detection pipeline from Kay et al. [37] – we train a YOLOv5 model for 5 epochs with the longest side of an image set to 896 and no augmentations. We remove duplicate predictions using non-maximal suppression. We use mAP50 [22] to evaluate detection performance frame-by-frame. We use a pretrained-frozen ByteTrack tracker and calculate MOTA [6], HOTA [46], and IDF1 [54] scores for evaluation. More details and hyperparameter settings are in the Supplemental Materials Sec. B.3 and B.4. For counting, we use trajectories from the tracks to create nMAE scores, defined in Kay et al. [37], for each domain. The tracking and counting pipelines do not require training.
For BUV, we follow the training procedure of Lin et al. [44]; we also follow their final fine-tuning step and evaluation to generate an AP50 metric. Note that we know that breast lesions are darker spots in ultrasounds. As a result, we invert our reconstruction error to take the minimum positive difference rather than the maximum:
$\text{inv}S_{t,T}=\min(0,I_{t}-I_{t-T})+I_{t}+\min(0,I_{t}+I_{t+t})$ (6)
Denoising for Classification. For POCUS, we perform 5-fold cross-validation as in Born et al. [8, 7], ensuring that frames from the same video are all in the same fold. We use the fine-tuning strategy and hyperparameters from Born et al. [8, 7]. We calculate each class’s precision, recall, and F1 scores and then average the folds’ metrics to determine overall metrics.
5 Results
We find that SAVeD is able to accurately denoise objects of interest in low signal-to-noise video. It improves a range of downstream tasks in a way that is computationally less resource-intensive and yields higher performance than other denoising methods.
5.1 Denoising Performance.
SAVeD produces clear contiguous objects, where other methods do not, shown in Fig. 3 & 4.
Figure 4: Qualitative denoising performance on CFC22. We can see that the fish is easiest to spot as a bright patch after processing with our denoiser. The green box highlights the fish location. Each denoised image zooms in to that green bounding box. The red arrow in the raw frame points to the fish location. Additional example visualizations are in Fig. 14 in Supp. Mat.
Fish Denoising: CFC22. SAVeD increases the contrast between fish and background, see Fig. 4. As such, the distributions of pixel intensities at the same location when fish are present and when they are not are distinct. This is shown in Tab. 1, where SAVeD’s $PSNR_{D_{KL}}$ is significantly higher than that of other methods.
Avg. KL-Divergence ($\uparrow$   )
Train Val Test
Raw 1005 652 860
CFC22++[37] 63.7 368 458
Median-filtered[27] 523 334 364
Gaussian-filtered[27] 793 507 756
N2V[40] 227 194 180
UDVD[59] 402 254 272
Denoised (framewise) 494 405 548
SAVeD (Ours) 1366 994 1458
Table 1: $PSNR_{D_{KL}}$ Quantitative denoised methods KL-divergence metric between P(Fish) vs. Q (Non-fish) as distributions of pixel intensities. For all ground truth bounding boxes, P and Q are composed as follows: P – we take the set of pixels in each box from frames with objects. Q – we extract the set of pixels from the same box location from a frame where there is no object at that location. Raw=raw noisy frame $I_{t}$, CFC22++[37] = 3-channel image (raw, background-subtracted, frame-to-frame difference), Denoised(framewise)= denoised with $[\Phi,\Omega,\Psi]$ trained with $I_{t}$ as the target (i.e. no motion augmentation), SAVeD=denoised with motion augmentation, as in Sec. 3.2. We calculate the KL-divergence metric, discussed in Sec. 3.4.1. $\uparrow$ indicates the metric is better the larger it is. Best values are bolded, worst values are in italics.
Fluorescent Cells Denoising: Fluo.  We found that our denoising method increases the cells’ brightness relative to the background, as seen in Fig. 3. As is standard[59], and because the data size is small, we only perform qualitative analysis on Fluo.
5.2 Detection Performance
SAVeD outperforms other denoising methods when evaluated on downstream detection tasks of CFC22 and BUV.
Fish Detection: CFC22. The detection performance of SAVeD denoised frames is better than detection performance of other denoised frames for CFC22. This is shown in Fig. 4 & 5 and Tab. 2. SAVeD improves detection performance in areas where objects and signal are rare. Our denoised frames result in an improvement of 43.2% and 9.4% test accuracy compared to the raw and background-subtracted frames respectively, and a 5.1% boost in performance compared to a three-channel image (raw, background-subtracted, and frame-to-frame-absolute difference) described as baseline++ in Kay et al. [37], but hereon referred to as CFC22++.
CFC22 (Test) POCUS (5-fold-CV) BUV(Test)
[…]
Classical
Baseline 73.8 37.4 54.8 82.6 82.0 80.4 46.4
Median-Filter[27] 73.7 37.8 53.0 86.2 85.5 85.3 52.4
Mean-Filter[27] 76.4 44.3 41.4 84.0 84.7 83.2 52.6
Gaussian-Filter[27] 74.9 27.6 56.8 84.1 84.3 83.3 46.5
Blind-Spot/Mask Networks
N2V[40] 67.2 34.2 34.3 83.7 82.7 82.4 46.6
UDVD[59] 67.2 28.1 41.9 83.7 84.6 83.4 49.9
DAEs
AE 67.8 34.3 41.7 82.3 84.7 82.1 46.9
UNet [56] 73.9 34.1 56.3 83.7 84.6 83.4 51.0
UNet3D[56] 66.9 32.4 35.4 83.5 80.1 80.6 47.6
SAVeD (Ours) 77.6 47.4 33.9 87.5 86.7 86.3 59.5
Table 2: Downstream results. SAVeD does well across all datasets and downstream tasks. Best performance is bolded. Baseline refers to raw for medical ultrasound (POCUS [7, 8] and BUV [44]) and the strengthened baseline CFC22++[37] for fish sonar (CFC22 [37]. AP=average precision, AR=average recall, F1=average F1, mAP50=mean average precision of detections at IOU threshold 0.5, MOTA=Multi-Object Tracking Accuracy[6], nMAE=normalized mean absolute counting error[37]. More tracking results are in Fig. 6.
Figure 5: Denoising improves detections where signal is infrequent. (a) the ground truth fish patch locations (from bounding box labels) normalized over the dataset; most fish pass by in the top region, fish crossings below are infrequent, thus there is more training signal in the top part of the videos. (b/c) patchwise detection performance of CFC22++[37] and SAVeD, repsectively, on the CFC22 dataset. Heatmaps indicate mAP50 performance over all frames of the test set at pixel patches. The more red a patch is, the higher the mAP50 of that patch; the more blue the patch is, the lower the mAP50. (d) the difference, SAVeD - CFC22++, with solid ellipses at regions of heightened performance and dashed ellipses around areas of lowered performance. Denoising improves detections in areas where signal is infrequent. On the other hand, detection performance declines in areas where signal is abundant. Additional patch maps can be seen in Fig. 8 in the Supp Mat.
Compared to the background-subtracted frames, there is a 10.5% reduction in error in the validation set and a 20.3% reduction in error on the test set. SAVeD reduces error by 5.76% and 14.5% compared to the CFC22++ frames on the validation set and test set respectively.
Breast Lesion Detection: BUV. SAVeD clarifies the breast lesion imagery, as seen in Fig. 3. As a result, it is significantly more accurate than other denoising methods on breast lesion detection. This is shown in Tab. 2.
5.3 Tracking and Counting Performance
Fish Tracking and Counting: CFC22. Compared to classical and other DNN-based denoising methods, frames denoised by SAVeD achieve higher downstream performance for tracking and counting. The distinction between fish and background in the denoised frames is stronger, leading to fewer false negatives and more true positives. Results can be seen in Fig. 6 and Tab. 2.
Figure 6: Quantitative tracking improvements through denoising. CFC22++ consists of the three-channel (background-subtracted, absolute-difference, raw) frames. The “Perfect Denoiser” refers to frames that have black backgrounds and white masks at the bounding box locations. Denoising results in higher MOTA scores for val and test; SAVeD boosts IDF1 and HOTA scores in test moreso than in val.
5.4 Categorization Performance
Lung Health Categorization: POCUS. Our method yields the best 5-fold cross-validation image classification score compared to classical and network-based denoising methods on lung categorization, shown in Tab. 2. Fig. 7 shows the precision-recall for the denoising methods on the Covid class – SAVeD has the highest accuracy for Covid classification. Additional per-class performance analysis is in Sec. A.2 of the Supplementary materials.
Figure 7: Covid Precision-Recall across denoising methods. SAVeD has the highest average precision and average recall across denoising methods. Additional class-wise performance comparisons are in Fig. 10 in the Supplementary materials.
5.5 Ablations
We ablate the reconstruction target and the denoising autoencoder to find their relative importance.
Reconstruction Target. PFDwT1 is the most effective reconstruction target for increasing the accuracy of downstream tasks. We compared PFDwT1 to PFDwT2, $\sigma$ (the standard deviation over input frames), $\Sigma-5\bar{I}$ (the sum of 5 consecutive frames - 5*mean frame), and $\Sigma-3\bar{I}$ (the sum of 3 consecutive frames - 3*mean frame). The results are shown in Tab. 3 and Tab. 5(d) in the Supplementary material.
Autoencoder vs. no Autoencoder. Using a DAE improves downstream detection performance over using the reconstruction targets alone for all targets. This can be seen in Tab. 3 and Tab. 5(d) in the Supplementary material.
Architectures Our small denoising architecture has better performance on downstream tasks compared to larger architectures. Comparisons of SAVeD with a vanilla Autoencoder, a UNet[56], and a UNet[56] with 3D convolutional kernels are shown in Tab. 2. For additional details and architectures, see Tab. 5 in the Supp. mat. The vanilla Autoencoder’s architecture is also explicitly defined in Tab. 7.
mAP50
Signal Modification AE Train Val Test
Signal Modification w/o Denoising Network
Raw ($I_{t}$) 79.6 69.6 54.2
$\sigma$ 79.8 69.4 72.5
$\Sigma-5\bar{I}$ 78.3 67.6 71.7
PFDwT1 80.2 66.9 68.2
PFDwT2 81.2 68.1 63.0
Signal Modification w/ Denoising Network
Raw ($I_{t}$) 81.5 68.4 73.4
$\sigma$ 82.2 70.0 73.5
$\Sigma-5\bar{I}$ 79.8 68.1 71.7
PFDwT1 83.5 70.6 77.6
PFDwT2 82.2 68.5 71.4
Table 3: Effect of Different Motion Enhancements with and without the Denoising Network on CFC22. All detectors that leverage the DAE have superior performance to those that use only the motion-enhanced target on the test set. The modified signal is used as the reconstruction target for the denoising autencoder when it is present, and is the input signal for the downstream task when the autoencoder is not used. All results are on CNNs with skip connections with resolution 1024 and bottleneck 512.
6 Discussion
While SAVeD is a clear improvement, we recognize that there are further improvements to be made and research directions to be explored.
Limitations. As the spatiotemporal component of our method relies an object’s location to be overlapping in sequential frames, very fast moving objects may decrease performance on downstream tasks using SAVeD. On the other hand, if objects are stationary, SAVeD does not improve performance, though it also should not be detrimental.
Future work. We are interested in training end-to-end: combining the representations from the denoiser and the downstream tasks. We are also interested in experimenting with more than 3 frames as input to broaden the motion signature. Finally, we recognize the shared qualities of each of these datasets and also understand that self-supervised methods are data-hungry [43, 61]. As such, one could explore the performance benefit of training on all datasets collectively to learn general low-SNR video properties.
7 Conclusion
We present SAVeD, a self-supervised denoising method that does not require noise-free video which improves downstream performance in low-SNR videos. This is based on the confirmed intuition that while there is motion in the foreground and background, their motion signatures are distinct, and a simple model can separate them to improve the signal-to-noise ratio. Our proposed method captures objects’ motion while leveraging autoencoders’ denoising capabilities to improve downstream task performance efficiently. Our approach is general and applicable to a range of low-SNR video tasks and domains.
Supplementary Material
We present additional experimental results as ablations Sec. (A), additional implementation details (Sec. B), and additional visualizations (Sec. C).
Benefits and risks of this technology. Improving classification, tracking, and counting in sonar and ultrasound videos is useful across medical, ecological, and other fields. Counting fish with sonar allows for a non-invasive way to measure population size, which can then be used for conservation and ecological efforts, for understanding effects of climate change, and for monitoring human fishing behavior for economical reasons. Improving classification in ultrasound videos, too, paves a path for more automated diagnosis. Risks, though, are inherent in both tracking applications and applications of sensitive data. Care must be taken when using these models, so that they are not used blindly without human intervention to make decisions.
Appendix A Additional Experimental Results
A.1 Additional CFC22 Ablation Results
As in the main paper, we evaluate CFC22 on the detection val/test splits, and show results using mAP50 across the dataset splits. We look at the effect of bottleneck size in the hourglass network, traditional augmentations, input resolution size, and reconstruction targets on how the trained denoiser affects downstream detection performance.
Bottlenecks size. For all experiments on CFC22, we use a default input size of 1024hx512w, reconstruction target as PFDwT1, mean-squared error (MSE) loss, and we train the denoiser for 20 epochs. Here the hourglass network remains 2 layers, with the number of input channels as 512, but the number of channels in the middle layer changes. We notice that for training, larger (less-restrictive) bottlenecks yield higher performance. For val and test, though, bottleneck sizes over 64 improve performance, but the differences between 128 and 512 is worse for val and negligible for test. Results can be seen in Tab. 5(a).
Resolution size. We vary the input resolution size to train the denoiser and notice higher performance for train and test when higher resolutions are used, seen in Tab. 5(b). We hypothesized that higher resolution size would make the denoiser more stable for downstream detections because higher resolution sizes would mean that removing entire fish (i.e. small fish) would be less probable. It is interesting to note that the highest resolution size 2048x1024 for val led to lower detection performance than that of resolution size 1024x512. We note, though, that higher resolutions lead to smaller batch sizes and longer training time.
Traditional Augmentations. We apply salt-and-pepper noise, gaussian-blur, motion-blur, brightness, and erasing from the kornia.Augmentations library. We found that no traditional augmentations, though, improve downstream detection performance. Results can be seen in Tab. 5(c).
Reconstruction Targets. We experimented with a handful of reconstruction targets:
Frame difference—such as absolute difference ($S_{|d|}=|I_{t}-I_{t+T}|$) or raw difference ($S_{d}=I_{t}-I_{t+T}$)—has been used in other self-supervised works as a spatiotemporal reconstruction target [62]. This works well in video where the movement in the background is less than the foreground movement. For our experiments, we use absolute difference as frame difference.
Raw frame ($I_{t}$) predicts the input frame alone.
Background subtraction (bs) We approximate the background frame, $\bar{I_{v}}$, as the mean aggregate of video over time. This is based on the approximation that objects of interest are sparse in terms of space and time. The mean frame is subtracted from every frame in the video ($S_{d}=(I_{v})_{t}-\bar{I_{v}})$.
Positive Frame Difference with current frame (PFDwTN). We discuss this in section 3.2. We experimented with T=2 (PFDwT2) and T=1 (PFDwT1), ultimately selecting T=1.
Standard Deviation across all frames (sigma) is taken across all of the frames loaded in a window of continuous frames, $\sigma(I_{t-N}:I_{t+N})$ where 2N+1 is the size of the window. We experimented with N=1 and N=2.
Sum frames minus N*background ($\Sigma-N\bar{I}$) sums all of the frames in a window size N and takes the positive difference $N*\bar{I}$ where $\bar{I}$ is the mean frame of all frames in a video: $\max(0,(\sum_{t}^{T}I_{t})-N\bar{I})$. We experimented with window sizes N=3 and N=5.
Visualizations of all of these can be seen in Fig. 9
CFC22 POCUS BUV
N2V[40] 12 days 0.75 hours 1.5 hours
UDVD[59] 8 days* 12 hours 23 hours
SAVeD 20 hours 0.5 hours 2 hour
Table 4: SAVeD is time-efficient. Note that UDVD took 8 days* to train CFC22, but UDVD trained CFC22 only for one epoch. For all other datasets, UDVD trained for 10 epochs all on 2 NVIDIA RTX 4090 GPUs.
(a) One clip from the CFC22-train river. You can can see the trajectory and patchwise detection performance improves after denoising. Overall, the biggest denoising gains appear to be at the edges of the cone, where fish are known to be small (entering/exiting) but moving.
(b) One clip from the CFC22-val river. The denoising gain is smaller and therefore more difficult to see here.
(c) One clip from the CFC22-test river.
Figure 8: Denoising-improved detection leads to better tracks. On the single-clip trajectory plots, orange dots indicate false negatives, green dots indicate true positives, red indicates false positives.
Figure 9: Reconstruction Targets. The window T=5 set of frames is shown with each reconstruction target we experimented with on CFC22. While $\Sigma-N\bar{I}$ frames appear strong in this example, we found that empirically they struggled to capture fish that did not move significantly between frames.
Train Val Test1
Bottleneck mAP50 mAP50 mAP50
64 79.1 68.6 71.6
128 80.0 69.2 72.6
512 81.6 69.4 72.6
(a) Bottleneck size. A larger bottleneck outperforms overly-constricted networks. All results are from CNNs with no skip connections and non-residual blocks.
Train Val Test1
Resolution mAP50 mAP50 mAP50
512 81.3 69.2 71.9
1024 79.1 68.6 71.6
2048 80.1 68.1 72.1
(b) Resolution size. There is no clear optimal - in terms of train and val, the smallest resolution size is the best; however, in terms of test, the largest resolution size is optimal. Note that higher resolutions also lead to longer training times.
Train Val Test1
Augmentations mAP50 mAP50 mAP50
$\text{saltpepper}_{0.25}$ 81.2 68.5 72.2
$\text{saltpepper}_{0.5}$ 83.7 69.7 75.1
$\text{saltpepper}_{0.75}$ 81.4 69.2 72.8
$\text{gaussianblur}_{0.25}$ 82.1 69.9 74.8
$\text{gaussianblur}_{0.5}$ 81.3 68.9 75.0
$\text{gaussianblur}_{0.75}$ 83.5 68.4 75.6
$\text{motionblur}_{0.25}$ 83.5 68.3 76.5
$\text{motionblur}_{0.5}$ 81.2 68.2 74.7
$\text{motionblur}_{0.75}$ 83.7 69.6 73.9
$\text{brightness}_{0.25}$ 83.7 69.8 74.7
$\text{brightness}_{0.5}$ 82.2 69.0 73.9
$\text{brightness}_{0.75}$ 83.6 69.7 76.8
$\text{erase}_{0.25}$ 82.0 68.7 68.0
$\text{erase}_{0.5}$ 81.1 68.7 75.6
$\text{erase}_{0.75}$ 77.4 59.3 62.4
(c) Augmentations. Augmentations appear to degrade performace. All augmentation experiments are named as $augmentation_{probability}$.
Train Val Test1
Target mAP50 mAP50 mAP50
Raw* 81.5 68.4 73.4
Absolute Difference $|I_{t}-I_{t+1}|$ 81.6 69.2 73.5
Sigma(N=5) 78.8 69.2 72.8
$\hat{S}_{t,T=1}$* 82.7 70.0 74.0
$\hat{S}_{t,T=2}$* 82.8 70.6 73.0
$\hat{S}_{t,T=2}+\hat{S}_{t,T=1}-I_{t}$* 83.7 69.2 74.6
$\Sigma-3\bar{I}$ 80.3 68.3 69.0
$\Sigma-5\bar{I}$ 80.7 68.7 72.0
(d) Reconstruction targets. Reconstruction targets including both the original frame and the next or previous frames do better than reconstruction targets incorporating information from just one. Reconstruction targets with the current frame in have *. All results are on CNNs with resolution 1024 and bottleneck 512 (with no SKIP connection).
Train Val Test1
Architectures mAP50 mAP50 mAP50
Autoencoder 82.6 68.9 67.8
CNN-fine 82.7 69.1 74.0
CNN-SKIP 83.5 70.6 77.6
CNN-residual 83.5 69.2 73.1
CNN-resnet-block 79.8 70.0 73.6
UNet-downscaled 82.1 69.1 75.8
UNet 81.2 70.0 73.9
UNet3D 79.0 67.0 66.9
(e) Denoising backbone architecture. All experiments have our target from equation 2 ($\hat{S}_{t,T=1}$) as their target. Networks are ordered from smallest (in terms of parameters and TFLOPs) to largest – it is interesting to note that as model size increases, performance does not necessarily increase. We see the top performer is the CNN-SKIP architecture.
Table 5: Additional denoise-detection ablations on CFC22. All values are generated via the detection stage of our pipeline. All reconstruction targets are sized 1024 x 512 unless otherwise stated. We report the mAP50 of the combined background-subtracted and target reconstruction frame unless otherwise noted. Default settings are marked in gray.
A.2 POCUS Per-Class Performance
SAVeD performs well across all classes (COVID, Pneumonia, and Regular) in the POCUS dataset (Fig. 10). For Pneumonia, precision levels across all methods were lower than for other classes. Pneumonia false negatives are more often categorized as Regular than they are Covid across all denoising methods.
Figure 10: SAVeD (starred) has high precision and high recall across all POCUS classes.
Appendix B Implementation Details
B.1 SAVeD Architecture Details
Our method uses a series of convolution blocks with skip connections as an encoder $\Phi$, a bottleneck (hourglass network) $\Theta$, and a reconstruction decoder $\Psi$. Architectural details about each of these are shown in Tab. 6. For more implementation details, the code will be made publicly available.
Encoder
Type Input shape Output shape
Conv_block (1,1024,512) (16, 1024, 512)
Pooling (16, 1024, 512) (16, 512, 256)
Skip (16, 1024, 512) (16, 512, 256)
Conv_block (16, 512, 256) (32, 512, 256)
Pooling (32, 512, 256) (32, 256, 128)
Skip (32, 512, 256) (32, 256, 128)
Conv_block (32, 256, 128) (64, 156, 128)
Pooling (64, 156, 128) (64, 128, 64)
Skip (64, 156, 128) (64, 128, 64)
Conv_block (64, 128, 64) (128, 128, 64)
Pooling (128, 128, 64) (128, 64, 32)
Skip (128, 128, 64) (128, 64, 32)
Conv_block (128, 64, 32) (256, 64, 32)
Pooling (256, 64, 32) (256, 32, 16)
Skip (256, 64, 32) (256, 32, 16)
Conv_block (256, 32, 16) (512, 32, 16)
Pooling (512, 32, 16) (512, 16, 8)
Skip (512, 32, 16) (512, 16, 8)
Decoder
Type Input shape Output shape
Upsample_block (512, 16, 8) (256, 32, 16)
Skip_connect (256, 32, 16) (768, 32, 16)
Conv_block (768, 32, 16) (512, 32, 16)
Upsample_block (512, 32, 16) (256, 64, 32)
Skip_connect (256, 64, 32) (512, 64, 32)
Conv_block (512, 64, 32) (256, 64, 32)
Upsample_block (256, 64, 32) (128, 128, 64)
Skip_connect (128, 128, 64) (256, 128, 64)
Conv_block (256, 128, 64) (128, 128, 64)
Upsample_block (128, 128, 64) (64, 256, 128)
Skip_connect (64, 256, 128) (128, 256, 128)
Conv_block (128, 256, 128) (64, 256, 128)
Upsample_block (64, 256, 128) (32, 512, 256)
Skip_connect (32, 512, 256) (64, 512, 256)
Conv_block (64, 512, 256) (32, 512, 256)
Upsample_block (32, 512, 256) (1, 1025, 512)
Table 6: Architecture details of the encoder, bottleneck, and decoder of SAVeD.Conv_block" is a basic convolutional block composed of 3x3 convolution with padding side of 1 and ReLU activation. “Skip" is a skip connection (stored to be input into the decoder) composed by maxpooling and then running a 1x1 convolution. “Upsample_block" is a 2D ConvTranspose with a 2x2 kernel and a stride of 2 and a ReLU activation. “Skip_connect” is the concatenation of the output from Upsample_block+Conv_block and the “Skip” corresponding to the same layer saved by the encoder. Note that this architecture is on input size of 1024x512.
Encoder
Type Input shape Output shape
Conv_block (3, 1024, 512) (16, 1024, 512)
Pooling (16, 1024, 512) (16, 512, 256)
Conv_block (16, 512, 256) (32, 512, 256)
Pooling (32, 512, 256) (32, 256, 128)
Conv_block (32, 256, 128) (64, 156, 128)
Pooling (64, 156, 128) (64, 128, 64)
Conv_block (64, 128, 64) (128, 128, 64)
Pooling (128, 128, 64) (128, 64, 32)
Conv_block (128, 64, 32) (256, 64, 32)
Pooling (256, 64, 32) (256, 32, 16)
Conv_block (256, 32, 16) (512, 32, 16)
Pooling (512, 32, 16) (512, 16, 8)
Decoder
Type Input shape Output shape
Bilinear_upsample_block (512, 16, 8) (512, 32, 16)
Conv_block (512, 32, 16) (256, 32, 16)
Bilinear_upsample_block (256, 32, 16) (256, 64, 32)
Conv_block (256, 64, 32) (128, 64, 32)
Bilinear_upsample_block (128, 64, 32) (128, 128, 64)
Conv_block (128, 128, 64) (64, 128, 64)
Bilinear_upsample_block (64, 128, 64) (64, 256, 128)
Conv_block (64, 256, 128) (32, 256, 128)
Bilinear_upsample_block (32, 256, 128) (32, 512, 256)
Conv_block (32, 512, 256) (16, 512, 256)
Bilinear_upsample_block (16, 512, 256) (16, 1024, 512)
Conv_block (16, 1024, 512) (1, 1024, 512)
Table 7: Architecture details of the vanilla autoencoder.Conv_block" is a basic convolutional block composed of 3x3 convolution with padding side of 1 and ReLU activation. “Bilinear_upsample_block" is a Bilinear Upsample kernel with a scale factor of 2 and align corners set to True. Note that this architecture is on input size of 1024x512.
B.2 SAVeD Hyperparameters and Time Comparisons
The hyperparameters for our method are in Tab. 8. All DAE models are trained until the training loss converges on 2 NVIDIA RTX 4090 GPUs. Tab. 4 shows how much time each denoiser took on each dataset.
Dataset Resolution Target Epochs Batch size Learning Rate Optimizer Scheduler
CFC22 (1024,512) $S_{t,T=1}$ 20 16 0.0005 AdamW Plateau f=0.1 pat=2
POCUS (1024,512) $S_{t,T=1}$ 120 8 0.0005 AdamW Step ss=2, $\gamma=0.05$
BUV (1024,1024) ${\text{inverse}(S_{t,T=1})}$ 40 8 0.0005 AdamW Step ss=2, $\gamma=0.05$
Fluo (1024,1024) $I_{t}$ 1000 8 0.0005 AdamW Step ss=2, $\gamma=0.05$
Table 8: SAVeD Hyperparameters. Note “$\text{inverse}(S_{t,T=1})$$=\min(0,I_{t}-I_{t-T})+I_{t}+\min(0,I_{t}-I_{t+T})$. f=Factor, pat=Patience, ss=Step size.
Figure 11: Visualization of $PSNR_{D_{KL}}$. Both images on the left are noisy images. The image on the far left has a fish located in the red bounding box. The image in the middle is a frame from the same video clip but with no fish in the red box. The histogram compares the pixel intensity values of the pixels within the bounding boxes. We can see these distributions, while overlapping, are distinct.
B.3 CFC22 Detector Details
We fine-tune a YoloV5-small model pretrained on COCO using the default training settings from Ultralytics over 5 epochs with a batch size of 16. As in Kay et al. [37], we resize all inputs to have 896 pixels as their longest side; the learning rate is 0.0025. We select the best model checkpoint based on validation mAP50. We train on two NVIDIA RTX A6000 GPUs. We recognize that the number of epochs (5) differs from the number of epochs in the original paper (150), and that is two-fold: 1.) CFC22++ Val and Test Performance after 5 epochs are $<1\%$ lower than Val and Test Performance after 150 epochs, therefore our denoised improvement beats the CFC22++ method also after CFC22++ is trained for 150 epochs while the detector model based on SAVeD frames is trained for 5 epochs; 2.) We wanted to show that a very simple detector could be used as a result of passing in denoised frames.
Figure 12: Denoising lowers detections error-rates by improving precision and recall (a) shows baseline detection error (1-mAP50) compared to our detection error after our denoising preprocessing step. For all splits train, val, and test, denoising results in lower error. (b) compares error rates from the validation set (x-axis) to error rates from the test set (y-axis) to see how denoising impacts each split. There is a 5.8% reduction in error in the val set and a 14.5% reduction in error on the test set. (c) Shows inverted Precision-Recall plots for each CFC22 dataset split – precision and recall both improve for all splits.
Figure 13: Breakdown of track performance improvements for CFC22 val and test. We can see test improves far more than val, as is standard for the CFC22 dataset.
B.4 CFC22 Tracker Details
We use a pretrained ByteTrack tracker with hyperparameters selected as the optimal hyperparameters for tracking performance on the validation set. Max age, the time until a missing or occluded object is assigned a new id, is 20; Min hits, the minimum number of frames with a track for the track to be considered valid, is 11; IOU threshold, the iou required for an object to be considered the same in the subsequent frame, is 0.01.
Appendix C Visualizations
Additional visualizations of the denoising performance on fish in sonar (CFC22[37]) can be seen in Fig. 14).
Figure 14: Additional visualizations of denoising methods on CFC22